### v1

# Direct Neural Machine Translation with Task-level Mixture of Experts models

Abstract
Direct neural machine translation (direct NMT) is a type of NMT system that translates text between two non-English languages. Direct NMT systems often face limitations due to the scarcity of parallel data between non-English language pairs. Several approaches have been proposed to address this limitation, such as multilingual NMT and pivot NMT (translation between two languages via English). Task-level Mixture of expert models (Task-level MoE), an inference-efficient variation of Transformer-based models, has shown promising NMT performance for a large number of language pairs. In Task-level MoE, different language groups can use different routing strategies to optimize cross-lingual learning and inference speed. In this work, we examine Task-level MoE’s applicability in direct NMT and propose a series of high-performing training and evaluation configurations, through which Task-level MoE-based direct NMT systems outperform bilingual and pivot-based models for a large number of low and high-resource direct pairs, and translation directions. Our Task-level MoE with 16 experts outperforms bilingual NMT, Pivot NMT models for 7 language pairs, while pivot-based models still performed better in 9 pairs and directions.
1 Introduction
Figure 1: Task-level MoE model, with LP - based routing; each Language Pair is routed through a top-2 router to an expert in the model experts’ layer. From a pretrained Task-level MoE model we can extract a smaller dense network specializing in a certain task, e.g. ja-ko Neural Machine Translation.
Neural machine translation (NMT) (Sutskever et al., 2014; Bahdanau et al., 2014; Firat et al., 2016a; Johnson et al., 2017; Zoph and Knight, 2016; Bahdanau et al., 2014) has made remarkable progress in recent years, particularly with the advent of the Transformer architecture (Vaswani et al., 2017), which has shown impressive results in bilingual and multilingual translation tasks (Aharoni et al., 2019; Hassan et al., 2018; Tang et al., 2021). Large Multilingual Neural Machine Translation models in particular have shown noteworthy performance and can serve translations between any languages, even for language pairs not seen in training, for which though NMT quality is often still low. The focus so far has primarily been on English-centric NMT, as there is an abundance of English-centric MT data (Tiedemann, 2018) but few parallel data for non-English (low-resource) language pairs. Hence the need for improving methods for Direct NMT, i.e. NMT between non-English languages has become evident (Yang et al., 2021; Zhang et al., 2020).
Direct NMT methods, when no xx–yy data is available (where xx represents the source language code, and yy represents the target language code), can be categorized into Zero-shot and Zero-resource approaches. In Zero-shot NMT (Johnson et al., 2017; Gu et al., 2019; Wang et al., 2021; Philip et al., 2020), translation occurs as follows: either through a bridging language, typically English, where models are trained on En–yy and xx–En datasets without ever being exposed to xx–yy data, while evaluation takes place on the few existing xx–yy pairs; or evaluating large multilingual models on xx–yy pairs, with neither xx or yy seen during model training. Zero-resource NMT (Currey and Heafield, 2019; Firat et al., 2016b; Bapna et al., 2022), on the other hand, generates pseudo-parallel data between direct language pairs, utilizing these synthetic xx–yy datasets during training. Yet bilingual or pivot-based models for direct NMT, though showing impressive NMT quality, are not practical and cost-efficient, especially as the number of languages for translation increases. Additionally, relying on related languages when evaluating a multilingual model on an unseen pair does not guarantee good NMT performance. At the same time, scaling up to multilingual models of billions of parameters, trained on very large datasets of multiple language pairs, induces large training and inference costs.
Sparse models, and specifically, sparsely activated Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Lepikhin et al., 2020; Zoph et al., 2022; Ryabinin and Gusev, 2020), shape a promising direction in creating a universal translation model between all languages, while addressing efficiency concerns. We are the first to investigate performance of a specific type of MoE models, Task-level MoE models (kudugunta-etal-2021-beyond-distillation) in the context of Direct Neural Machine Translation (NMT), focusing on the impact of different task-to-expert mapping methods on model performance and expert utilization. In our experiments, we train large multilingual Task-level MoE models across various configurations, including Zero-shot, Zero-resource, and scenarios where direct language pair data is available. We conduct a comparative analysis of BLEU scores across Task-level MoE models with different mapping strategies and expert counts, as well as bilingual NMT and pivot-based NMT models; our findings serve as a valuable resource for determining the optimal model configuration to achieve high-quality translation results in various direct language pairs. By training large sparse multilingual Task-level MoE models which perform well for direct pairs, and showcase training efficiency, we may be able to extract smaller expert-specific dense models, which specialize in translating between Language Pairs or translating into one particular language. These models are ready for deployment and can replace bilingual or pivot-based models (Figure 1). Prior work on Task-level MoEs kudugunta-etal-2021-beyond-distillation focused on demonstrating the desired inference efficiency of task-level routing, compared to other routing approaches (token- or sentence-level), for the task of Multilingual Neural Machine Translation (MNMT) between English-centric pairs (En-yy and xx–En). To the best of our knowledge, we’re the first ones to examine performance of Task-level MoE models in Direct NMT.
ja–th bg–mk ru–tr fr–ar $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ Task-level MoE 21.51 25.36 26.44 23.29 10.24 8.72 19.37 15.06 Bilingual 6.24 7.26 24.47 21.28 9.53 7.58 19.6 9.71 Pivot-level 18.34 39.28 28.24 27.03 18.44 13.84 21.57 14.91 Number of Experts 16 64 16 16 16 16 16 16 Decoding mapping method lp_b tl_b tl_a tl_a tl_a tl_a tl_a tl_a
zh–ko ja–vi ja–ko ja–zh $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ Task-level MoE 34.35 28.70 25.17 19.18 42.64 41.38 34.94 22.52 Bilingual 5.71 3.16 8.20 19.6 6.12 4.47 3.16 7.26 Pivot-level 28.13 35.24 18.97 19.61 21.49 36.69 19.98 27.46 Number of Experts 16 16 16 64 16 16 16 64 Decoding mapping method lp_b lp_b lp_b lp_a lp_b lp_b lp_b lp_a
Table 1: Best BLEU scores from our Task-level MoE models trained with 16 or 64 experts, for direct pairs. We compare to Bilingual and Pivot-level NMT models’ scores and state the Task-level MoE model/setup under which best scores were obtained. overall best scores per pair are highlighted. lp_x decoding mapping method implies a model trained with Language Pair (LP) - based task to expert mapping, and tl_x implies target language (TL) - based task to expert mapping.
ja–th bg–mk ru–tr fr–ar $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ Task-level MoE (LP) lp_a 0.47 2.26 3.21 22.75 1.55 0.74 - - lp_b 21.51 18.92 3.07 20.43 4.86 0.83 19.37 15.06 lp_c 0.42 1.50 5.97 18.32 5.57 0.80 2.40 0.86 Task-level MoE (TL) tl_a 1.15 0.78 26.44 23.29 10.24 8.72 5.36 1.18 tl_b 1.39 6.09 5.19 10.15 4.45 0.86 5.62 0.81 Bilingual 6.24 7.26 24.47 21.28 9.53 7.58 19.60 9.71 Pivot-level 18.34 39.28 28.24 27.03 18.44 13.84 21.57 14.91 Number of train sentences 6,200 205,651 1,467,160 17,163,359
zh–ko ja–vi ja–ko ja–zh $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ Task-level MoE (LP) lp_a 2.81 4.09 0.77 5.32 3.33 11.78 1.67 9.06 lp_b 34.35 28.70 25.17 1.78 42.64 41.38 34.93 2.50 lp_c 2.35 3.04 0.57 1.31 2.71 10.14 1.30 2.26 Task-level MoE (TL) tl_a 4.51 0.61 0.87 14.31 3.87 1.30 1.86 18.77 tl_b 4.44 10.7 1.34 4.80 3.91 19.91 2.79 4.06 Bilingual 5.71 3.16 8.20 19.6 6.12 4.47 3.16 7.26 Pivot-level 28.13 35.24 18.97 19.61 21.49 36.69 19.98 27.46 Number of train sentences 498,968 604,940 974,896 1,339,622
Table 2: BLEU scores of Task-level MoE models trained with 16 experts, for direct pairs, in both directions, for models with Language Pair (LP) - or target language (TL) - based routing during training, lp_a, lp_b, lp_c and tl_a, tl_b mapping during inference, respectively. We compare to Bilingual and Pivot-level NMT models’ scores and highlight the model/setup under which best scores were obtained.
2 Related Work
Direct Neural Machine Translation
Firat et al. (2016b) first proposed a multilingual multi-way zero-resource NMT model which outperformed previous pivot-based approaches, laying the foundation for numerous subsequent zero-shot and/or zero-resource NMT methods. Various researchers have since expanded upon this foundational work, exploring different techniques and strategies for improving direct NMT (Ji et al., 2020; Johnson et al., 2017; Zhang et al., 2020; Chen et al., 2017; Lu et al., 2018; Al-Shedivat and Parikh, 2019; Rios et al., 2020; Zheng et al., 2017; Gu et al., 2019; Arivazhagan et al., 2019; Kim et al., 2019; Cheng and Cheng, 2019; Chen et al., 2018; Currey and Heafield, 2019; Ballesteros and Sanderson, 2003; Lakew et al., 2021). More specifically, Freitag and Firat (2020) constructed a multi-way aligned training set from existing training corpora, aligning examples in different language pairs with identical source or target sentences. They introduced a hierarchical data sampling strategy to prevent the over-representation of English in the training set and employed a temperature-based strategy to select target languages, while choosing source languages uniformly. Their complete Multilingual Neural Machine Translation models demonstrated improved performance on direct pairs and comparable to baseline results on English-centric pairs. Liu et al. (2020) discretized the encoder output latent space into entries in a learned codebook, enabling the representation of source sentences in a common language and increasing model robustness and performance in zero-shot setups. More recently, eriguchi-etal-2022-building pre-trained and fine-tuned a many-to-many multilingual NMT model that surpassed several pivoting and bilingual baselines for Direct NMT, and ElNokrashy et al. (2022), presented a novel approach in which source and target language tokens are appended to the encoder input, while target language tokens are added to the decoder input. This differs from previous methods that used Source-only tokens and beginning of Sentence tokens for the encoder and decoder inputs, respectively, and aims to improve performance of Direct NMT models by providing more relevant context to both the encoder and decoder.
Latest works include the approach of Yang et al. (2022), who unified three distinct models to guide the student model during Direct NMT, aiming to improve its performance by harnessing the strengths of multiple models; and Xu et al. (2022), who proposed a Direct NMT method comprising two stages: extraction of highly similar examples and their translations across different languages to construct a multi-way aligned parallel dataset, which enables the model to learn from a richer set of translation examples, and generation of additional aligned examples using a well-trained generative model.
bg–mk fr–ar ja–th zh–ko ja–vi ja–ko ja–zh $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ Task-level MoE (LP) lp_a 19.18 - - 16.33 17.50 21.34 19.08 19.18 25.3 25.25 22.52 lp_b 19.07 4.47 2.13 0.44 2.79 2.05 2.51 0.47 0.96 4.46 0.71 lp_c 19.04 2.69 1.18 0.36 0.50 0.59 0.47 0.45 0.80 2.05 0.56 Task-level MoE (TL) tl_a 20.29 7.37 5.38 8.30 25.33 12.61 20.27 5.95 14.25 26.39 0.01 tl_b 20.35 7.32 5.50 8.33 25.36 12.43 20.34 5.96 14.26 26.40 10.49 Bilingual 21.28 9.71 19.60 6.24 7.26 5.71 3.16 8.20 6.12 4.47 3.16 Pivot-level 27.03 14.91 21.57 18.34 39.28 28.13 35.24 18.97 21.49 36.69 19.98 Number of train sentences 205,651 17,163,359 6,200 498,968 604,940 974,896 1,339,622
Table 3: BLEU scores of Task-level MoE models trained with 64 experts, for direct pairs, in both directions, for models with Language Pair (LP) - or target language (TL) - based routing during training, lp_a, lp_b, lp_c and tl_a, tl_b mapping during inference, respectively. We compare to Bilingual and Pivot-level NMT models’ scores and highlight the model/setup under which best scores were obtained.
Conditional Computation, Sparsity and Mixture-of-Experts models
Fedus et al. (2022) extensively review sparse Mixture-of-Experts (MoE) models in Deep Learning, starting from the foundational work of Shazeer et al. (2017), Lepikhin et al. (2020) and Fedus et al. (2021). Their work delves into the scaling characteristics, routing algorithms, and training improvements proposed for MoE models, as well as the latest applications of Sparse MoE models across various domains, such as Natural Language Processing (NLP) (Zoph et al., 2022; Chowdhery et al., 2022; Lee-Thorp and Ainslie, 2022; Du et al., 2022), Computer Vision (Riquelme et al., 2021; Puigcerver et al., 2020; Wu et al., 2022; Hwang et al., 2022), Speech Recognition (You et al., 2021, 2022), and Multimodal Learning (Mustafa et al., 2022). In NLP, numerous studies have explored the potential of Sparse MoE models for diverse applications. Kudugunta et al. (2021) introduced the concept of task-level routing in MoE models, achieving improved translation scores compared to large multilingual models with the same inference cost and dense student models distilled from token- or position-level MoE models. By using task-level MoE models in our experiments, we investigate and optimize their performance in direct pairs’ NMT.
3 Experiments
Model
We train sparse encoder-decoder Task - level MoE models, with dim = 1024, hidden dim = 4096, 8 heads, 3 layers. Following kudugunta-etal-2021-beyond-distillation, we similarly replace the Transformer Feed Forward Network (FFN) with a set of identical FFN experts. We use a 32,000 token SentencePiece vocabulary (Kudo and Richardson, 2018), shared in both the encoder and decoder of the model. We experiment with the number of experts, selecting either 16 or 64 experts, which correspond to models with approximately 1 billion and 3.5 billion parameters, respectively. We investigate the impact of different task id to expert mapping methods, setting the task id to be either the Target Language or the Language Pair for translation. Following Kudo and Richardson (2018); Kudugunta et al. (2021), data sampling temperature is T=5 during training, and we train 8 models for 2 million steps.
The training process has a batch size of 256, a maximum sentence length of 128. By examining performance of these models under different configurations, we want to better understand the relationship between the number of experts, task id to experts’ mappings, and translation quality, ultimately informing the optimal design choices for Task-level MoE models in various scenarios.
Languages and Training Datasets
We train our models using in-house production-scale datasets, consisting of English-centric, and Direct pairs’ parallel sentences, covering 108 languages, including English. Our training data includes a total of 107 English-centric pairs, and 53 direct pairs, with data sizes for each language pair ranging from a few thousand to a few billion sentences for each pair. The number of sentences in each parallel pair in the train set is provided in Table 4 in the Appendix.
Evaluation Sets and Task Mapping
For evaluation, we use internal test sets from two distinct sources: the Web Domain and Wikipedia. These test sets vary in size, containing anywhere from 500 to 5,000 sentences for each language pair. In both training and evaluation datasets, we follow a specific preprocessing convention, prepending the source sentence with <4xx> <2yy> tokens. These tokens serve as a cue for the model, helping it identify the source and target languages in context, and the model to better learn and adapt to various language pair combinations.
Our models use either the exact train set Language Pair (LP) or the target language (TL) of the pair as the task id for mapping translation pairs to different experts. In the former case, we directly map the specific language pair (En–yy, xx–En or xx–yy) to a task id, and have a total of 214 tasks (107 x 2 English-centric pairs) when training with English-only datasets, or 267 tasks (+ 53 direct Language Pairs) when incorporating Direct LPs. On the other hand, when using the target language as the task id, we have 108 tasks (107 languages and English) in all scenarios. There we leverage and try to benefit from the shared characteristics of translations into the same target language. Since we evaluate our models on direct (non-English) data Language Pairs, it is often the case that the exact mapping of the pair to an expert was not defined during training. To address this issue, we experiment with several approaches for mapping language pair xx–yy to a task id at inference time - in each method mentioned below we name the component used as task id:
  • Language Pair (LP) as task id.
    • lp_mapping_a (lp_a): exact LP
    • lp_mapping_b (lp_b): En–yy LP
    • lp_mapping_c (lp_c): xx–En LP
  • Target language as task id
    • tl_mapping_a (tl_a): yy
    • tl_mapping_b (tl_b): xx
4 Results & Discussion
Figure 2: Difference between BLEU scores of Task-level MoE models trained with 16 and 64 experts, for direct pairs, for models with Language Pair (LP) - based routing during training, lp_a, lp_b, lp_c mapping during inference, respectively, and models with Target Language (TL) - based routing during training, tl_a, tl_b mapping during inference. For each pair, we mark the BLEU score value of the model which shows the largest difference among the 16 and 64 experts Task-level MoE variation.
In this study, we have two primary baselines for comparison: bilingual NMT models and pivot-based NMT models, which are trained using a bridging strategy through English. We use BLEU (Papineni et al., 2002) for evaluating the translation quality of all direct language pairs. We are aware that learned metrics like COMET Rei et al. (2022) and BLEURT Sellam et al. (2020) show higher correlation with human judgements for high resource, English centric language pairs Freitag et al. (2022). However, it is unknown how well learned metrics perform on low resource languages and thus we decide to stick with BLEU for the main part of the paper. Tables 2 and 3 summarize results for Task-level MoE models with either 16 or 64 experts.
In Table 1, we show the specific configurations that led to the highest BLEU scores for Task-level MoE models - comparing those to the performance of bilingual and pivot-based models, we aim to emphasize the best BLEU scores for each language pair. This analysis allows us to identify the most effective models for each pair and gain insights into how different model configurations and expert counts contribute to translation performance, so we can better understand the advantages and limitations of Task-level MoE models with varying numbers of experts, as well as bilingual and pivot-based NMT models. Our results ultimately help inform the choice of translation models and strategies to optimize performance for different language pairs
Performance of NMT models varies significantly depending on the pair, and on the direction of translation - certain pairs show large discrepancies between translation directions. We see that pivot-based models perform better than bilingual ones for the majority of pairs. Bilingual NMT and pivot-based NMT models demonstrate strong results in several instances, while Task-level MoE models show mixed performance across different Language Pairs.
Which model performs best for each pair?
Table 1 results reveal that for the majority of direct language pairs, the highest Task-level MoE BLEU scores are achieved using a model with 16 experts, with the tl_a and lp_b mapping-to-experts methods surpassing other approaches. It is particularly noteworthy that for pairs of languages belonging to closely related language families (Japanese, Chinese, Korean, Vietnamese, and Thai), the LP-mapping-based models seem to perform best. In contrast, TL-based models tend to outperform the rest of the other language pairs. In this case, Task-level MoE models show superior performance in the forward translation direction across all pairs. Meanwhile, pivot-based models yield the best results in the backward translation direction for four out of five pairs in those related languages. This observation suggests that different mapping-to-experts strategies may be more effective for certain language pairs or translation directions, depending on the linguistic relationships and unique characteristics of the languages involved; a tailored approach in employing Task-level MoE models and mapping-to-experts methods is the key to optimizing translation performance for different language pairs and directions.
Comparison of different task to expert mapping methods
From Table 2, pivot-based models, Task-level MoE (LP) models with lp_b task id mapping and Task-level MoE (TL) models with tl_a task id mapping (both trained with 16 experts) are primarily the best performing models among all pairs.
Task-level MoE models with LP task id to expert routing and lp_b mapping at inference time excel in translations to Japanese, from Thai, Vietnamese, Korean, and Chinese (+3.17, +6.2, +21.15, +14.95 BLEU points over the pivot-level NMT model, respectively, and +15.27, +16.97, +36.52, +31.77 BLEU gain against bilingual models) as well as Korean-Chinese (ko-zh) (+6.22 BLEU compared to pivot-based model, +28.64 BLEU over the bilingual baseline) and French-Arabic (fr–ar) (+0.15 and +5.35 BLEU over the pivot and bilingual NMT models, respectively). This indicates the lp_b mapping strategy is particularly effective for these language pairs; routing to the expert best at En–yy translation, where yy is each pair’s target language, instead of routing to the expert explicitly trained on the exact direct pair, xx–yy, likely captures better the linguistic nuances of the pair and leads to improved translation scores.
Task-level MoE models trained with the TL task-to-expert mapping approach and the tl_a task id mapping, which leverages target language pairs for routing during inference, achieve the highest scores for translations involving Kazakh (kk), such as Kazakh-Russian (kk–ru) in both directions (+13.62 and +13.68 BLEU gain against pivot-based model, +0.34 and +12.34 BLEU points over bilingual models). Success of tl_a mapping strategy suggests the model shows superior performance in those cases and when the pair is properly routed to the expert specialised in the pair’s target language.
Table 3 shows results from our evaluation of 64 experts’ Task-level MoE models. Interestingly, pivot-based methods stand out as top performing in the majority of pairs and directions. These pairs include Bulgarian–Macedonian (bg–mk), French–Arabic (fr–ar), and Japanese–Thai (ja–th), as well as Chinese-Korean (zh-ko) in both directions, and Japanese-Korean (ja–ko). In all other pairs and directions, specifically when translating to Japanese from Vietnamese, Korean, and Chinese, Task-level MoE (LP) models with the lp_a mapping approach emerge as the best-performing models (+0.21, +3.81, +2.54 BLEU points against the pivot-based model, respectively, and +10, +19.18, +19.36 BLEU gain over the bilingual baseline).
For certain Language Pairs with smaller amounts of training data bilingual NMT models achieve reasonable performance even with limited data and often outperform others. On the other hand, we see that pivot-based NMT models perform well when there is a substantial amount of training data (e.g., Belarusian–Macedonian, and Russian–Turkish in both directions, Arabic–French, Chinese–Korean, Japanese–{Thai, Vietnamese, Chinese}). This might be due to the fact that these models leverage English as an intermediate bridging language between source and target languages, which is helpful when there’s enough data to help in learning meaningful representations.
Comparison of models with different number of experts
In Figure 2, we present the difference in BLEU scores on select direct language pairs for Task-level MoE models trained using Language Pair (LP) or Target Language (TL) based mapping, with 16 and 64 experts. We seek to visually emphasize differences in performance across models that utilize varying numbers of experts, in order to get insights into their relative effectiveness and understand how the choice of mapping strategy (LP or TL) and the number of experts (16 or 64) impacts translation performance. For the majority of pairs (Vietnamese–Japanese, and both directions of French–Arabic, Japanese–Thai, Chinese–Korean, Japanese–Korean, Japanese–Chinese), Task-level MoE models with 16 experts outperform those with 64. For those language pairs, with the exception of Japanese–Chinese, for which TL-based Task-level MoE with tl_a shows best performance, LP-based Task-level MoE with lp_b mapping show the largest gain when trained with fewer experts, corroborating our previous findings. For other pairs, such as Japanese–Thai, Japanese–Chinese, and Bengali–Macedonian, different model variations show a largest improvement when trained with either 16 or 64 experts. We also notice a significant, yet not maximum, BLEU score gain for a large number of pairs with the TL-based Task-level MoE model and the LP-based Task-level MoE model, when using tl_a and lp_a on the task-to-expert mapping during inference, respectively, between 16 and 64 experts’ models.
Routing decisions analysis
In Figures 3, 4, 5, 6 we show the experts’ utilization for the last model’s encoder and decoder layers for different pairs, for our TL-based Task-level MoE model, with tl_a used as a task to expert mapping approach during training; The darker a cell is in each of the Figures, the more that expert is used by the specific language pair. We can observe that language pairs with the same target language get properly routed to the same expert, as expected. It’s interesting to see that there are no major differences between the encoder’s routing decisions across checkpoints, yet we notice changes in experts’ assignment in the decoder during and at the end of training; at first, pairs with the same target language are mapped to the same different experts, yet as the model converges, we see a significant overlap in the selected experts, for pairs with different target languages. We also see a great overlap in experts chosen in the encoder during and at the end of training - the majority of experts are the same throughout model training and certain experts are preferred over others. This preference differs from the experts chosen in the decoder, where we see almost no overlap. It is also worth noting that the number of common experts between the encoder and the decoder is minimal, both during training and upon model convergence.
In the Appendix, we additionally show experts’ utilization in the encoder and decoder of our 64 experts’ TL-based Task-level MoE model.
Figure 3: Routing decisions of the last encoder layer of our Task-level MoE model with 16 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 1M steps.
Figure 4: Routing decisions of the last decoder layer of our Task-level MoE model with 16 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 1M steps.
Figure 5: Routing decisions of the last encoder layer of our Task-level MoE model with 16 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 2M steps.
Figure 6: Routing decisions of the last decoder layer of our Task-level MoE model with 16 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 2M steps.
5 Conclusions
We conducted a thorough analysis of the usage of Task-level Mixture of Expert models for Direct Neural Machine Translation. Our experiments reveal the strengths and weaknesses of different approaches and shed light on which configurations work best for specific direct language pairs. Our comparisons help in identifying best-performing models and offer valuable insights into how varying numbers of experts, and different task-to-expert mapping methods, during training and inference, can influence direct pairs’ translation quality in Task-level MoE models. Specifically, we noticed that Task-level MoE NMT models, along with pivot-based approaches, are frequently top performers for numerous direct language pairs. However, their performance varies between different pairs and translation directions. Besides observing NMT quality results, our analysis of expert’s usage throughout training also serves as an informative way to visualize and understand mapping of language pairs to model experts in the encoder and decoder throughout training.
Limitations
Training and evaluation of Task-level MoE models can be very challenging due to the models’ size and complexity. There is a large number of parameters that require tuning, such as the number and size of experts, the number and diversity of languages, and of English-centric and direct Language Pairs in training, the batch size, the vocabulary size and the maximum sentence length. At the same time, it is time and computationally expensive to evaluate on a large number of direct pairs, train separate Bilingual and Pivot-level models for all to use as baselines, and also perform the experts’ analysis and visualizations for those pairs. This automatically restricts the amount and breadth of results, and calls for further exploration of the model capabilities in the future.
Ethics Statement
Working with large language models raises several ethical concerns in NLP research, particularly related to quality, bias and toxicity of the output result Bender et al. (2021); Chowdhery et al. (2022); Blodgett et al. (2020); Brown et al. (2020). In the context of NMT, involvement of many stakeholders in the task, as well as the dangers arising from mistranslation of the original text need to be taken into consideration, since they can potentially affect the perception of the work and harm the interests of the different parts (Taivalkoski-Shilov, 2019; Gambier and Van Doorslaer, 2016). Undoubtedly, the benefits of responsibly developed and deployed NMT systems lie in making an author’s work more accessible, enabling the transfer of ideas to other audiences, and enhancing both the reader’s capabilities and the translator’s role and sense of contribution (Besacier, 2014), which are directions any work in the area should aim to focus on.
Appendix
In Figures 7, 8 we see the routing decisions maps for the last encoder and decoder layer of TL- based task-level MoE models trained with 64 experts and tl_a mapping during inference. The distribution to the experts is similarly expected here, following the training and inference pairs-to-experts’ mapping, as same target language pairs get routed to the same experts. The overlap in the encoder and decoder experts is minimal. All but one selected experts are different between the encoder and decoder.
Figure 7: Routing decisions of the last encoder layer of our Task-level MoE model with 64 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 2M steps.
Figure 8: Routing decisions of the last decoder layer of our Task-level MoE model with 64 experts, trained with pair target language to task id mapping, with tl_a used during inference, for 2M steps.
LP Num. of sentences | LP Num. of sentences | LP Num. of sentences
ar–fr 0 fr–ar 0 be–ru 3,641
ru–be 3,641 ja–th 6,200 th–ja 6,200
ko–th 6,208 th–zh 12,305 zh–th 12,305
id–ms 24,535 kk–tr 69,476 tr–$k$ 69,476
ug–en 129,688 tk–en 140,247 en–tk 144,094
en–ug 167,741 kk–ru 177,931 ru–$k$ 177,931
bg–mk 205,651 mk–bg 205,651 or–en 262,099
cs–es 285,546 es–cs 285,546 ko–vi 369,594
vi–ko 369,594 et–fi 466,254 fi–et 466,254
ko–zh 498,968 zh–ko 498,968 vi–zh 504,037
zh–vi 504,037 en–$t$ 556,751 de–ru 585,161
ru–de 585,161 ja–vi 604,940 vi–ja 604,940
zh–tr 609,066 en–or 658,015 en–rw 792,559
tt–en 819,579 cs–fr 895,912 fr–cs 895,912
ja–ko 974,896 ko–ja 974,896 cs–de 1,007,448
de–cs 1,007,448 rw–en 1,264,813 ja–zh 1,339,622
zh–ja 1,339,622 ru–tr 1,467,160 tr–ru 1,467,160
cs–ru 1,644,795 ru–cs 1,644,795 en–la 1,809,003
en–xh 1,895,974 la–en 2,167,039 en–sm 2,353,608
de–es 2,886,854 de–fr 2,886,854 es–de 2,886,854
en–st 2,931,225 en–sn 3,022,722 en–ig 3,058,540
en–haw 3,393,044 en–yo 3,404,254 en–lo 3,573,485
fr–de 3,681,694 ru–uk 3,792,978 uk–ru 3,792,978
en–yi 3,904,022 en–ku 5,254,260 en–ny 5,408,239
en–so 5,472,253 xh–en 5,728,449 sm–en 5,730,016
lo–en 5,766,691 en–mi 5,793,624 en–co 6,847,178
ig–en 7,224,372 en–tg 8,136,182 sn–en 8,319,539
yi–en 8,347,482 en–zu 8,376,755 haw–en 8,646,782
en–ha 8,995,196 en–sd 9,095,582 en–ps 9,446,484
yo–en 9,565,981 st–en 10,052,460 en–ky 10,213,875
mi–en 10,383,802 en–ht 11,758,532 ku–en 11,984,479
en–ga 12,494,184 ny–en 12,616,421 so–en 12,745,144
en–su 14,088,043 en–am 14,583,643 en–mt 15,693,771
en–pa 15,784,955 en–ceb 15,965,215 ga–en 16,145,438
en–mg 17,474,049 co–en 17,490,449 en–be 17,645,900
en–eo 18,104,746 es–ru 18,363,087 ru–es 18,363,087
tg–en 18,509,169 en–fy 19,495,193 ha–en 19,687,379
en–mn 20,012,644 sd–en 20,337,227 es–fr 20,792,781
fr–es 20,792,781 zu–en 21,010,498 ky–en 21,357,639
ps–en 21,613,753 ht–en 22,455,117 en–gd 22,495,036
fr–ru 22,890,960 ru–fr 22,890,960 en–eu 23,279,562
en–my 23,848,386 en–hy 25,779,512 en–lb 25,943,545
be–en 28,585,822 en–hmn 29,258,771 pa–en 29,592,573
su–en 30,234,526 en–jv 30,315,732 en–uz 30,943,271
en–si 31,847,809 en–$k$ 32,075,645 en–cy 32,089,017
am–en 32,110,524 en–ml 32,335,094 mg–en 32,356,401
en–gu 32,742,711 fy–en 33,389,347 en–mk 33,546,951
en–mr 33,629,504 hy–en 34,146,553 eu–en 35,158,779
hmn–en 35,795,085 en–bs 35,801,135 gd–en 36,082,889
en–kn 36,472,587 en–km 37,229,149 en–hr 37,248,537
en–ne 38,606,074 en–sw 41,198,888 mt–en 41,723,852
LP Num. of sentences | LP Num. of sentences | LP Num. of sentences
en–ka 42,992,730 en–te 43,112,561 en–gl 45,053,738
lb–en 46,042,055 kn–en 46,847,179 en–sr 48,786,193
mn–en 49,510,283 cy–en 50,252,795 en–af 50,525,196
gu–en 53,217,408 en–sq 55,374,368 my–en 55,886,059
jv–en 57,400,218 ceb–en 59,055,133 mk–en 61,867,654
en–ta 62,352,671 eo–en 62,552,455 en–az 64,274,455
kk–en 64,724,987 ne–en 69,288,148 si–en 69,667,030
en–lv 69,998,174 ml–en 70,059,552 en–is 71,837,680
hr–en 72,301,373 en–ur 74,947,168 en–et 75,581,372
te–en 75,967,139 en–bn 80,362,347 km–en 80,514,558
en–uk 86,997,037 en–ca 87,267,893 en–sk 89,304,972
gl–en 89,695,691 bs–en 90,263,002 sw–en 90,493,834
af–en 94,492,215 en–fil 97,193,343 en–sl 97,924,827
ka–en 99,344,373 ta–en 100,659,244 en–lt 102,081,855
is–en 103,425,627 sq–en 104,573,595 mr–en 105,247,834
uz–en 105,701,543 az–en 115,304,502 en–iw 117,411,687
en–ms 117,562,320 en–fa 117,720,531 en–bg 128,893,255
en–fi 133,307,056 en–el 135,547,629 en–ro 136,723,562
ca–en 137,014,332 sr–en 137,913,360 en–hu 146,210,128
ur–en 149,852,215 fil–en 156,986,036 lv–en 163,533,602
en–no 165,068,512 en–cs 167,802,683 en–da 176,676,171
sl–en 189,776,212 bn–en 190,650,787 et–en 197,134,423
uk–en 209,888,151 en–sv 217,128,777 lt–en 245,092,644
bg–en 289,282,624 sk–en 295,623,418 fi–en 300,151,301
el–en 322,429,692 th–en 340,378,951 ro–en 352,986,741
en–th 353,001,311 fa–en 402,836,484 ms–en 414,693,796
iw–en 427,057,571 hu–en 436,082,952 hi–en 490,464,288
en–hi 494,638,634 ar–en 525,411,304 id–en 527,345,119
en–ar 532,778,336 ko–en 535,863,357 en–id 538,581,029
da–en 539,573,001 en–ko 546,191,335 no–en 613,221,604
vi–en 632,740,475 en–vi 656,115,977 cs–en 661,833,636
sv–en 800,882,060 ja–en 846,991,020 tr–en 850,154,583
en–tr 869,295,018 en–ja 876,842,917 it–en 998,195,505
en–it 100,981,5294 pl–en 111,935,3071 en–pl 1,141,598,628
pt–en 1,155,038,272 en–pt 1,184,401,180 en–zh 1,228,817,744
zh–en 1,238,691,743 ru–en 1,425,268,039 en–ru 1,455,103,126
nl–en 1,580,819,532 en–nl 1,587,530,791 de–en 1,680,270,443
en–de 1,695,095,726 en–fr 1,887,609,530 fr–en 1,922,463,803
en–es 2,419,825,975 es–en 2,435,228,645
Table 4: Number of Sentences of each Language Pair (LP) in our train set.